OpenStudy (anonymous):

-7=6x+1 what is the solution?

sammixboo (sammixboo):

First, subtract 1 on both sides of the equation, since you are originally adding it -7 = 6x + 1 -7 - 1 = 6x + 1 - 1 -7 - 1= 6x So what is -7 - 1?

OpenStudy (anonymous):

-8?

sammixboo (sammixboo):

Correct, so we have -8 = 6x Now we divide 6 on both sides of the equation -8/6 = 6x/6 -8/6 = x So that tells us that \(\dfrac{-8}{6}\) = x But we can simplify that fraction. Do you know how to simplify fractions?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

wouldnt it be -4/3?

sammixboo (sammixboo):

Yes =)

sammixboo (sammixboo):

-8/6 is simplified to -4/3

OpenStudy (anonymous):

But thats an improper fraction how do i change that to a mixed number?

sammixboo (sammixboo):

It is proper =) For it to be improper the denominator has to be greater than the numerator. In this case, it isn't

OpenStudy (anonymous):

But you see its not matching any of my awnser choices

sammixboo (sammixboo):

What are your choices?

OpenStudy (anonymous):

-1 7/8, 1 5/6, 6, 5 1/3 , -4 , -4/5, 1/7, -1 1/3, -4 1/2, 3 1/5 , -1/8, -1 3/5 , -3 1/3, 2 2/9 , 2 3/4, 4 1/9 , 0, 7 1/2.

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

good job sam ;)